OpenStudy (anonymous):

Solve: −3(x − 2) = 15 3 −3 −5.7 −7

OpenStudy (anonymous):

−3(x − 2) = 15 Distribute the -3 -3x + 6 = 15 Then subtract 6 from both sides. -6 -6 ---------------- -3x = 9 Then divide both sides by -3. /-3 /-3 ---------- x = -3

OpenStudy (anonymous):

\[−3(x − 2) = 15 \]divide by \(-3\) \[x-2=-5\]add 2 \[x=-3\]

OpenStudy (anonymous):

Thank you!